• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:05:16 on June 12, 2004

    Entry number 125406

    keyword=Beam Time Accounting
    
    	
    Hall A

    06/11/04 1600-2400
    Experiment #: E00-114
    Shift Leader: E. Burtin
    Shift Workers: Z.-E. Meziani, M. Mazouz

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.500.330.000.000.000.000.000.83
    Beam Available
    but Not in Use
    0.000.170.251.000.420.000.000.332.17
    Beam Not Available
    or Unacceptable
    0.000.000.420.000.000.000.000.000.42
    Accelerator Con-
    figuration Change
    1.000.330.000.000.581.001.000.674.58
    Experiment
    Ready
    1.000.830.750.000.581.001.000.675.83
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.170.251.000.420.000.000.332.17
    Hall not
    ready reasons
    Target0.000.170.250.000.000.000.000.330.75
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.001.000.420.000.000.001.42
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 17:20 end of energy modulation test Q1 trip 17:50 start beam current studies 5 and 8.8 muA 18:20 beam off 18:45 beam back 19:00 request access to reset Q1 20:30 access over, J. Benesh takes over 23:00 target warmup due to some not really understood reasons. (see entry in Halog)

    Data Entry by: Etienne Burtin